# Generalized Lanczos method for systematic optimization of neural-network quantum states

### v1

@@ II.5 The NQS Lanczos method
Finally, reinforcement learning is applied to optimize all amplitude networks $\text{ANet}_{j}$ (except $\text{ANet}_{0}$) of the superposition state $|\Psi\rangle$. This produces the final improved state $|\tilde{\Psi}\rangle$, on which the Hamiltonian expectation is calculated to obtain the final improved energy $\tilde{E}$.
@@ III Numerical results and discussions
where $J_{1}$ and $J_{2}$ are the coupling strengths of nearest-neighbor sites and the next nearest-neighbor sites (denoted as $\langle ij\rangle$ and $\langle\langle ij\rangle\rangle$), respectively. This model is strongly frustrated when $J_{2}/J_{1}$ is around $0.5$. It is challenging to theoretically study this region. As the ratio $J_{2}/J_{1}$ increases from 00, the system undergoes a phase transition [46] from a Neel antiferromagnet to a gapless quantum spin liquid, then to a valence-bond solid, and finally to a collinear antiferromagnet. In this section, we test the NQS Lanczos method in the highly frustrated region. The test results from different stages (SLL and RLL) of the method are presented to comprehensively demonstrate the effectiveness of the method.
In each test, $|\psi^{\mathrm{net}}_{0}\rangle$ is obtained through the optimization of the amplitude network with a sign structure, which is fixed to the Marshall sign rule [47, 35]. The following tests all use the checkerboard patterned sign structure. Additionally, due to the highly rugged landscape of the wave function, the optimization results can be influenced by several factors, such as the initialization method of the network parameters, the seed of the random numbers for the initialization, the learning rate, and the optimization algorithm. The networks to be trained in the supervised learning are initialized by Kaiming initialization [48]. To reduce the fluctuation caused by initialization, we use four different seeds to initialize the network and retain the result with the best optimization performance. The Adam algorithm is used for all optimization. All calculations in this work are performed on a single NVIDIA 3090 GPU.
@@ III.1 Test of the supervised learning Lanczos algorithm on the $4\times 4$ lattice
We first test our method on the $L=4$ square lattice with $J_{2}/J_{1}=0.55$. The test results are shown in Fig. 4. The improved energies are obtained from the SLL algorithm. The horizontal axis represents the number of the Lanczos steps. The vertical axis indicates the error of the improved energies with respect to the energy obtained from the exact diagonalization (ED) method. The error decreases exponentially from $6.32\times 10^{-3}$ at $p=0$ to $1.73\times 10^{-8}$ at $p=5$.
Figure 4: Improved energies on the $4\times 4$ lattice. The horizontal axis represents the number of Lanczos steps. The vertical axis indicates the error of the improved energies with respect to the energy obtained from ED. As the number of the Lanczos steps increases, the energy rapidly converges to the exact one.
For the $L=4$ square lattice, the Hilbert space can be traversed easily, the coefficients $a_{i}$ and $b_{i+1}$ can be calculated accurately, and the loss can be optimized to $10^{-8}$. In this case, the supervised trained states are almost strictly orthogonal. The combination of the Lanczos method, supervised learning, and NQSs is successful. The SLL algorithm can perfectly implement the Lanczos method without the need for the RLL algorithm.
@@ III.2 Test of the supervised-learning Lanczos algorithm on the $6\times 6$ and $10\times 10$ lattices
The performance of the SLL algorithm on the square lattices with $L=6$ and $L=10$ is tested in this section. For the $L=6$ lattice, the aCNN is used with the ResNet-v2 connection. We test the cases of $p=1$ and $p=2$, and the results are shown in Fig. 5. The horizontal axis represents the number of the Lanczos steps. The vertical axis represents the error ($(E-E_{\mathrm{ED}})\times 10^{4}$) of the improved energies with respect to the energy from ED. For $L=10$, we continue to use the aCNN with the ResNet-v1 connection and test the case of $p=1$, and the results are shown in Fig. 6. The points with $p=0$ indicate the energies of the initial Lanczos states, while for $p>0$, the points labeled by SLL indicate the improved energies obtained by the SLL algorithm.
Figure 5: Improved energies on the $6\times 6$ lattice. The model is test with $J_{2}/J_{1}=0.5$, $0.55$, and $0.6$. The horizontal axis represents the number of the Lanczos steps. The vertical axis indicates the error ($(E-E_{\mathrm{ED}})\times 10^{4}$) of the improved energies with respect to the energy obtained from ED. At $p\neq 0$, the points of SLL indicate the energies obtained by the SLL algorithm. At $p=0$, the points of SLL indicate the energies of the initial states. The points of RLL ($p=n$, $\text{ANet}_{j}$) indicate the energies obtained by the RLL algorithm. The superposition state consists of $n+1$ NQSs. $\text{ANet}_{j}$ indicates that the amplitude network of Lanczos step $j$ is optimized in the RLL algorithm.
@@
$L=6$ $|\psi^{\mathrm{net}}_{0}\rangle$ 95. 8% 96. 4% 98. 8% 0. 21 0. 15 0. 10
$|\psi^{\mathrm{net}}_{1}\rangle$ 86. 0% 87. 0% 87. 0% 0. 44 0. 44 0. 43
$L=10$ $|\psi^{\mathrm{net}}_{0}\rangle$ 91. 2% 89. 6% 0. 46 0. 32
@@
Table 2: Energies obtained from the NQS Lanczos method. SLL ($p=n$) represents the energy obtained by the SLL algorithm. RLL ($p=n$, $\text{ANet}_{j}$) denotes the energies obtained through the reinforcement learning of $\text{ANet}_{j}$, which is the amplitude network in the $j$-th NQS constituting the superposition state. The result of MinSR is from Ref. [29].
@@
RLL ($p=1$, $\text{ANet}_{1}$) -0. 503052(3) -0. 493142(5) -0. 488024(7)
RLL ($p=2$, $\text{ANet}_{1}$) -0. 503119(4) -0. 493357(5) -0. 488569(7)
RLL ($p=2$, $\text{ANet}_{1}$ $\text{ANet}_{2})$0. 503178(4) -0. 493471(5) -0. 488799(7)
@@
III.3 Test of the reinforcement-learning Lanczos algorithm on the $6\times 6$ and $10\times 10$ lattices
In the above tests, although the effectiveness of the combination of supervised learning with the Lanczos method has been demonstrated, the difficulty of loss convergence makes it impossible for the Lanczos state to be accurately represented by the NQS. In this section, we test the RLL algorithm on the superposition states obtained from the previous tests. For a superposition state $|\Psi\rangle$ that will be optimized by the RLL algorithm, all parameters of the sign networks and the superposition coefficients are fixed. The amplitude network $\text{ANet}_{j}$ ($j\neq 0$) will be optimized one by one until the energy no longer decreases.
@@ III.3 Test of the reinforcement-learning Lanczos algorithm on the $6\times 6$ and $10\times 10$ lattices
The results obtained by the RLL algorithm for the $L=6$ and $L=10$ square lattices are shown in Figs. 5 and 6, respectively. In these figures, RLL ($p=n$, $\text{ANet}_{j}$) denotes the energies obtained through the reinforcement learning for the superposition state, which consists of $n+1$ NQSs. $\text{ANet}_{j}$ indicate the $j$-th amplitude network is optimized. The specific energies are shown in Table 2 of Appendix B. The energy improvement after the reinforcement learning is substantial. For $L=6$, the improved energies achieved by optimizing only $\text{ANet}_{1}$ with $p=1$ are better than those from SLL ($p=2$). Moreover, the energy improvement increases further by optimizing $\text{ANet}_{1}$ and $\text{ANet}_{2}$ with $p=2$.
The above results show that the RLL algorithm is very effective for the adjustment of the amplitude. We speculate that this adjustment partially compensates for the insufficient amplitude optimization during the supervised learning and further improves the generalization capability of the network.
For the NQS Lanczos method, it might be sufficient to retain only those steps where significant energy improvement is observed, such as $p=1$ or $p=2$. Combining the SLL algorithm with the RLL algorithm, and using the optimized superposition state as the updated $|\psi^{\mathrm{net}}_{0}\rangle$, as shown in Fig. 3, is a promising path to further improve the results. Due to limited GPU computational resources, we have not done it yet. We will try it in the future.
@@ IV Summary and outlook
In summary, we integrate supervised learning, reinforcement learning, NQS, and the Lanczos method to develop a method, namely the NQS Lanczos method, to iteratively find the ground state of quantum many-body systems. It consists of two main innovations: the supervised-learning Lanczos (SLL) and the reinforcement-learning Lanczos (RLL) algorithms. This method is tested in high-frustration regions of the two-dimensional Heisenberg $J_{1}$-$J_{2}$ model on the square lattices with $L=4,6$, and $10$. The results demonstrate its effectiveness.
@@ Acknowledgements.
This work was supported by the National Key R&D Program of China (Grants No. 2024YFA1408601 and No. 2024YFA1408602), the National Natural Science Foundation of China (Grant No. 12434009), and the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0302402). Computational resources were provided by the Physical Laboratory of High Performance Computing in Renmin University of China.
@@ Appendix A Diagonalization of the Hamiltonian with a set of non-orthogonal basis states
$\{|\alpha_{0}\rangle,\cdots,|\alpha_{p}\rangle\}$ =$=\{|\psi^{\mathrm{net}}_{0}\rangle,\cdots,|\psi^{\mathrm{net}}_{p}\rangle\}S.$ (17)
@@ Appendix B Specific values in the test of the NQS Lanczos method
This section presents specific values in the test of the NQS Lanczos method on square lattices with $L=6$ and $10$. The energies obtained by the SLL algorithm and the RLL algorithm are listed in Table 2. The coefficients $a_{i}$ and $b^{2}_{i}$ of the SLL algorithm are listed in Table 3.

### v2

@@ II.6 The NQS Lanczos method
Finally, VMC is applied to optimize all amplitude networks $\text{ANet}_{j}$ (except $\text{ANet}_{0}$) of the superposition state $|\Psi\rangle$. This produces the final improved state $|\tilde{\Psi}\rangle$, on which the Hamiltonian expectation is calculated to obtain the final improved energy $\tilde{E}$.
@@ III Numerical results and discussions
where $J_{1}$ and $J_{2}$ are the coupling strengths of nearest-neighbor sites and the next nearest-neighbor sites (denoted as $\langle ij\rangle$ and $\langle\langle ij\rangle\rangle$), respectively. This model is strongly frustrated when $J_{2}/J_{1}$ is around $0.5$. It is challenging to theoretically study this region. As the ratio $J_{2}/J_{1}$ increases from 0, the system undergoes a phase transition [49] from a Neel antiferromagnet to a gapless quantum spin liquid, then to a valence-bond solid, and finally to a collinear antiferromagnet. In this section, we test the NQS Lanczos method in the highly frustrated region. The test results from different stages (SLL and VMC) of the method are presented to comprehensively demonstrate the effectiveness of the method.
To avoid confusion with traditional VMC, we introduce the term “VMC Lanczos” (denoted as VMCL) to specifically label the results obtained from performing VMC on the superposition state.
In each test, $|\psi^{\mathrm{net}}_{0}\rangle$ is obtained through the optimization of the amplitude network with a sign structure, which is fixed to the Marshall sign rule (MSR) [50, 35]. The following tests all use the checkerboard patterned sign structure. Additionally, due to the highly rugged landscape of the wave function, the optimization results can be influenced by several factors, such as the initialization method of the network parameters, the seed of the random numbers for the initialization, the learning rate, and the optimization algorithm. The networks to be trained in the supervised learning are initialized by Kaiming initialization [51]. To reduce the fluctuation caused by initialization, we use four different seeds to initialize the network and retain the result with the best optimization performance. The Adam algorithm is used for all optimization. All calculations in this work are performed on a single NVIDIA 3090 GPU.
@@ III.1 Test of the supervised learning Lanczos algorithm on the $4\times 4$ lattice
We first test our method on the $L=4$ square lattice with $J_{2}/J_{1}=0.55$. The test results are shown in Fig. 4. The improved energies are obtained from the SLL algorithm. The horizontal axis represents the number of the Lanczos steps. The vertical axis indicates the relative error of the improved energies with respect to the energy obtained from the exact diagonalization (ED) method, $\epsilon_{\text{rel}}=(E-E_{\text{ED}})/|E_{\text{ED}}|$. The relative error decreases exponentially from $1.20\times 10^{-2}$ at $p=0$ to $3.75\times 10^{-8}$ at $p=5$.
Figure 4: Improved energies on the $4\times 4$ lattice. The horizontal axis represents the number of Lanczos steps. The vertical axis indicates the relative error of the improved energies with respect to the energy obtained from ED, $\epsilon_{\text{rel}}=(E-E_{\text{ED}})/|E_{\text{ED}}|$.
As the number of the Lanczos steps increases, the energy rapidly converges to the exact one.
For the $L=4$ square lattice, the Hilbert space can be traversed easily, the coefficients $a_{i}$ and $b_{i+1}$ can be calculated accurately, and the loss can be optimized to $10^{-8}$. In this case, the supervised trained states are almost strictly orthogonal. The combination of the Lanczos method, supervised learning, and NQSs is successful. The SLL algorithm can perfectly implement the Lanczos method without the need for the VMC optimization part.
@@ III.2 Test of the supervised-learning Lanczos algorithm on the $6\times 6$ and $10\times 10$ lattices
The performance of the SLL algorithm on the square lattices with $L=6$ and $L=10$ is tested in this section. For the $L=6$ lattice, the aCNN is used with the ResNet-v2 connection. We test the cases of $p=1$ and $p=2$, and the results are shown in Fig. 5. The vertical axis indicates the rescaled relative error, $\epsilon_{\text{rel}}\times 10^{-3}$, of the improved energies with respect to the energy obtained from ED. For $L=10$, we continue to use the aCNN with the ResNet-v1 connection and test the case of $p=1$, and the results are shown in Fig. 6. The points with $p=0$ indicate the energies of the initial Lanczos states, while for $p>0$, the points labeled by SLL indicate the improved energies obtained by the SLL algorithm.
Figure 5: Improved energies on the $6\times 6$ lattice. The model is test with $J_{2}/J_{1}=0.5$, $0.55$, and $0.6$. The vertical axis indicates the rescaled relative error, $\epsilon_{\text{rel}}\times 10^{-3}$, of the improved energies with respect to the energy obtained from ED. At $p\neq 0$, the points of SLL indicate the energies obtained by the SLL algorithm. At $p=0$, the points of SLL indicate the energies of the initial states. The points of VMCL ($p=n$, $\text{ANet}_{j}$) indicate the energies obtained by the VMC. The superposition state consists of $n+1$ NQSs. $\text{ANet}_{j}$ indicates that the amplitude network of Lanczos step $j$ is optimized in the VMC.
@@
$L=6$ $|\psi^{\mathrm{net}}_{1}\rangle$ 95. 8% 96. 4% 98. 8% 0. 21 0. 15 0. 10
$|\psi^{\mathrm{net}}_{2}\rangle$ 86. 0% 87. 0% 87. 0% 0. 44 0. 44 0. 43
$L=10$ $|\psi^{\mathrm{net}}_{1}\rangle$ 91. 2% 89. 6% 0. 46 0. 32
@@
Table 2: Comparison of amplitude overlap for different loss functions on the $L=6$ square lattice.
   Loss    $J_{2}/J_{1}=0.5$    $J_{2}/J_{1}=0.6$
   MSE    0. 84626(3)    0. 92769(7)
   Fidelity s1    0. 82627(4)    0. 91683(6)
   Fidelity s2    0. 84742(3)    0. 91936(5)
To identify the most effective loss function for supervised learning, we compared their convergence behaviors. For each strategy, we tuned key hyperparameters, including the learning rate and its decay schedule. The results from eight random seeds were evaluated, with the best-performing instances detailed in Table 2. This table reports the overlap between the optimized neural network wave function and the target wave function ($p=1$) for the $L=6$ square lattice. The MSE loss demonstrates relatively superior performance across various reference points and benefits from a simple functional form, establishing it as the optimal choice for this study.
Exploring more effective optimization strategies for loss function based on fidelity or more suitable forms of the loss function remains highly valuable for future algorithmic improvements. For example, the $L^{2}$ distance discussed in Ref. [52] presents an alternative.
@@
Figure 8: Improved energies on the 1D $L=56$ chain with $J_{2}/J_{1}=0.5$. The vertical axis indicates the relative error of the improved energies with respect to the exact energy, $\epsilon_{\text{rel}}=(E-E_{\text{exact}})/|E_{\text{exact}}|$.
To convincingly demonstrate the utility of the SLL algorithm, we apply the SLL algorithm to the Heisenberg $J_{1}$-$J_{2}$ model on a 1D chain with $L=56$, which has an exact solution $E_{\text{exact}}=-0.375J_{1}$ for $J_{2}/J_{1}=0.5$. The results are shown in Figure 8. It can be observed that as $p$ increases, the relative error of the energy gradually decreases. Compared to $p=2$ ($\epsilon_{\text{rel}}=8.54\times 10^{-4}$), the error for $p=4$ ($\epsilon_{\text{rel}}=4.98\times 10^{-4}$) decreased by approximately half, indicating that the computations for $p=3$ and $4$ are effective. This suggests that as $p$ increases, the error continues to decrease.
Table 3: Sign accuracy of the NQS Lanczos method benchmarked against the exact solution on the $L=6$ square lattice, where the superposition state (with $p=2$) comprises three NQSs.
State $J_{2}/J_{1}=0.5$ $J_{2}/J_{1}=0.55$
initial (MSR) 0. 9804 0. 9352
superposition (SLL) 0. 9983(1) 0. 9960(1)
superposition (VMCL) 0. 9981(1) 0. 9955(1)
Table 4: Energies obtained from the NQS Lanczos method. SLL ($p=n$) represents the energy obtained by the SLL algorithm. VMCL ($p=n$, $\text{ANet}_{j}$) denotes the energies obtained by the optimization of $\text{ANet}_{j}$, which is the amplitude network in the $j$-th NQS constituting the superposition state. The result of MinSR is from Ref. [29].
@@
VMCL ($p=1$, $\text{ANet}_{1}$) -0. 503052(3) -0. 493142(5) -0. 488024(7)
VMCL ($p=2$, $\text{ANet}_{1}$) -0. 503119(4) -0. 493357(5) -0. 488569(7)
VMCL ($p=2$, $\text{ANet}_{1}$ & $\text{ANet}_{2})$ -0. 503178(4) -0. 493471(5) -0. 488799(7)
@@
III.3 Test of the VMC optimization part on the $6\times 6$ and $10\times 10$ lattices
In the above tests, although the effectiveness of the combination of supervised learning with the Lanczos method has been demonstrated, the difficulty of loss convergence makes it impossible for the Lanczos state to be accurately represented by the NQS. In this section, we test the VMC optimization part on the superposition states obtained from the previous tests. For a superposition state $|\Psi\rangle$ that will be optimized by the VMC, all parameters of the sign networks and the superposition coefficients are fixed. The amplitude network $\text{ANet}_{j}$ ($j\neq 0$) will be optimized one by one until the energy no longer decreases.
@@ III.3 Test of the VMC optimization part on the $6\times 6$ and $10\times 10$ lattices
The results obtained by the VMC for the $L=6$ and $L=10$ square lattices are shown in Figs. 5 and 6, respectively. In these figures, VMCL ($p=n$, $\text{ANet}_{j}$) denotes the energies obtained by the optimization of the superposition state, which consists of $n+1$ NQSs. $\text{ANet}_{j}$ indicate the $j$-th amplitude network is optimized. The specific energies are shown in Table 4 of Appendix B. The energy improvement after the VMC is substantial. For $L=6$, the improved energies achieved by optimizing only $\text{ANet}_{1}$ with $p=1$ are better than those from SLL ($p=2$). Moreover, the energy improvement increases further by optimizing $\text{ANet}_{1}$ and $\text{ANet}_{2}$ with $p=2$.
The above results show that the VMC optimization part is very effective for the adjustment of the amplitude. We speculate that this adjustment partially compensates for the insufficient amplitude optimization during the supervised learning and further improves the generalization capability of the network.
We further benchmarked the sign accuracy of the superposition states against the exact solution for the $L=6$ square lattice. The results are summarized in Table 3. The SLL algorithm significantly improves the sign accuracy from 98.04% ($J_{2}/J_{1}=0.5$) and 93.52% ($J_{2}/J_{1}=0.55$) of the initial states (with MSR imposed) to 99.83% and 99.60%, respectively, demonstrating its efficacy in refining the sign structure of initial states.
For the NQS Lanczos method, it might be sufficient to retain only those steps where significant energy improvement is observed, such as $p=1$ or $p=2$. Combining the SLL algorithm with the VMC optimization part, and using the optimized superposition state as the updated $|\psi^{\mathrm{net}}_{0}\rangle$, as shown in Fig. 3, is a promising path to further improve the results. Due to limited GPU computational resources, we have not done it yet. We will try it in the future.
@@ IV Summary and outlook
In summary, we integrate supervised learning, VMC, NQS, and the Lanczos method to develop a method, namely the NQS Lanczos method, to iteratively find the ground state of quantum many-body systems. It consists of two main parts: the supervised-learning Lanczos (SLL) algorithm and the VMC optimization part. This method is tested in high-frustration regions of the two-dimensional Heisenberg $J_{1}$-$J_{2}$ model on the square lattices with $L=4,6$, and $10$. The results demonstrate its effectiveness.
@@ Acknowledgements.
This work was supported by the National Key R&D Program of China (Grants No. 2024YFA1408602 and No. 2024YFA1408601) and the National Natural Science Foundation of China (Grant No. 12434009). Z.-Y.L. was also supported by the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0302402). Computational resources were provided by the Physical Laboratory of High Performance Computing in Renmin University of China.
Code availability
The code supporting the findings of this study is openly available in a GitHub repository via the URL provided in Ref.  [^1] The code supporting the findings of this study is openly available in a GitHub repository at https://github.com/QTMEC-RUC/nqs-lanczos (will be publicly accessible upon publication of the paper)..
@@ Appendix A Diagonalization of the Hamiltonian with a set of non-orthogonal basis states
$\{|\alpha_{0}\rangle,\cdots,|\alpha_{p}\rangle\}$ $=\{|\psi^{\mathrm{net}}_{0}\rangle,\cdots,|\psi^{\mathrm{net}}_{p}\rangle\}S.$ (19)
@@ Appendix B Specific values in the test of the NQS Lanczos method
This section presents specific values in the test of the NQS Lanczos method on square lattices with $L=6$ and $10$. The energies obtained by the SLL algorithm and the VMC optimization part are listed in Table 4. The coefficients $a_{i}$ and $b^{2}_{i}$ of the SLL algorithm are listed in Table 5.